# Language Model Preference Evaluation with Multiple Weak Evaluators

### v3

@@ 4 Theoretical Analysis
In this section, we provide a theoretical foundation for our method, showing that by modeling preference graphs as random perturbations of a ground truth DAG, GED can reliably recover the true structure through graph ensemble and denoising with high probability, demonstrating its robustness in handling noisy evaluations. Theoretically, we treat each of our preference graph as a random perturbation of some ground truth DAG $G=(V,A)$. Specifically, we consider a random graph generator $\mathcal{G}(G,\delta_{1},\delta_{2})$ with parameters $\delta_{1},\delta_{2}\in[0,1]$ such that $G_{i}=(V_{i},A_{i})\sim\mathcal{G}(G,\delta_{1},\delta_{2})$ satisfies $V_{i}=V$.
@@
Theorem 1.
[…]
where $G\subseteq\text{MAS}(\widehat{G})$ represents that $G$ is a subgraph of $\text{MAS}(\widehat{G})$ and $U=\frac{|V|(|V|-1)}{2}-|A|$ is the number of pairs of unconnected nodes in $G$.
The full proof is given in Appendix G. From the theorem, we can see that the probability of failure decreases exponentially as the number of samples $N$ increases. Meanwhile, this guarantee only requires $\delta_{1}<0.5$ and does not place restrictions on $\delta_{2}$, which are very mild conditions.
Table 1: Performance comparison of response selection methods across five benchmarks. GED consistently outperforms baseline methods, demonstrating the effectiveness of graph denoising and the aggregation of multiple evaluators.
@@
Multi-MV 66.18 29.57 74.77 86.42 11.72 53.73
GED (w/o denoising) 69.25 30.98 74.29 87.17 12.68 54.87
Multiple evaluator GED 70.73 32.44 75.58 88.18 13.33 56.05
Figure 2: Comparison of GED with GPT-3.5, GPT-4-o-mini, and GPT-4-o on 100 randomly selected tasks. GED consistently outperforms GPT-3.5 across all tasks and surpasses GPT-4-o-mini on challenging tasks like HumanEval and GSM8k, showcasing the effectiveness of weak evaluator aggregation with graph denoising.
5 Experiments on Response Selection
@@
In this section, we evaluate the performance of GED on five benchmarks: HumanEval (Chen et al., 2021), AlpacaEval (Li et al., 2023b), MATH (Hendrycks et al., 2021), GSM8k (Chen et al., 2021), and GAIA (Mialon et al., 2023). The Qwen2-72B (Yang et al., 2024a) model ($\mathcal{M}$) generates ten candidate responses per question, and we assess the effectiveness of different methods in selecting the best response. For further implementation details, see Appendix A. We evaluate performance using three setups. First, in the single model setting, the baselines include ContraSolver(Zhang et al., 2024b), Self-consistency(Wang et al., 2022), and direct evaluation with models (Llama3-8B, Mistral-7B, Qwen2-7B and Qwen2-72B). Additionally, we include a baseline called ListPreference, where instead of pairwise comparisons, all candidate responses are input into Qwen2-72B for selecting the most appropriate response. Then, in the single evaluator setting, individual evaluators (Llama3-8B, Mistral-7B, Qwen2-7B, Qwen2-72B) select the best response from $\mathcal{M}$’s outputs, with and without applying GED’s graph denoising. Finally, in the multiple evaluators setup, we combine three small evaluators (Llama3-8B, Qwen2-7B, Mistral-7B) to select responses from Qwen2-72B with GED. We also introduce a baseline, Multi-MV, which selects the response that receives the most votes from evaluators in pairwise comparisons. We present the results of GED and its variant (w/o denoising), which ensembles the preference graphs without the denoising step.
@@
GED (w/o ensemble) 62.82 68.44 68.44 69.34 67.34 67.27
GED (w/o denoising) 64.84 69.23 69.81 75.35 74.37 70.72
Multiple evaluator GED 66.59 71.14 71.14 77.17 76.46 72.50
@@
Table 1 presents the results of the response selection task across five benchmarks. GED consistently outperforms baseline methods, including both single model evaluations (single model) and direct response selection by individual models (single evaluator). This demonstrates the strength of aggregating weak evaluators with GED, particularly when coupled with graph denoising, which enhances response quality by filtering out noise and biases. Furthermore, by combining preference graphs derived from smaller models (Llama3-8B, Mistral-7B, Qwen2-7B), GED outperforms a much larger evaluator (Qwen2-72B). This underscores the value of ensemble methods in mitigating the limitations of individual evaluators.
Then, the denoising process proves to be crucial for improving consistency and overall response quality. The substantial performance gains observed when using GED with denoising, compared to both the single evaluator setup and the ensemble without denoising, highlight its importance in refining response selection. For Multi-MV, while it improves upon individual evaluators, it still underperforms GED, highlighting GED’s ability to capture nuanced evaluation signals and reduce inconsistencies. Additionally, we observed that the ListPreference baseline performed worse than Qwen2-72B as single evaluator, likely due to LLM limitations in handling long-text. Lastly, to further evaluate GED, we compared its performance with GPT-3.5, GPT-4-o-mini, and GPT-4-o. Due to computational and API cost constraints, we limited the evaluation to 100 data points for each task. As shown in Figure 2, GED consistently outperformed GPT-3.5 across all tasks and surpassed GPT-4-o-mini on challenging benchmarks like HumanEval and GSM8k. These results highlight the superiority of GED, particularly in leveraging multi-weak evaluators and graph denoising to outperform individual state-of-the-art models.
Ablation study.
We evaluate the impact of removing the ensembling step in GED, referred to as the (w/o ensemble) variant. In this case, individual evaluators’ preference graphs are denoised and converted to rankings, which are then aggregated using methods such as Weight Score, Kemeny, Weighted Kemeny, Pairwise Majority, and Weighted Pairwise Majority (detailed in Appendix N). For simplicity of presentation, we use Weight Score to represent GED (w/o ensemble) (Weight Score). As shown in Figure 3, all (w/o ensemble) methods consistently underperform compared to GED. This performance gap arises because converting graphs to ranks before aggregation leads to information loss. In contrast, GED ensembles the graphs directly, preserving more detailed preference information and resulting in better final rankings.
6 Experiments on Model Ranking
@@
In this section, we evaluate the effectiveness of GED in the model ranking task within a human preference setting, using the AlpacaEval benchmark (Li et al., 2023b). We employ 30 widely used models from the AlpacaEval dataset as our model set $\mathcal{M}$, while the benchmark’s questions form the question set $Q$. The rankings provided by the AlpacaEval benchmark serve as ground truth for evaluating the accuracy of various ranking methods. This is justified by AlpacaEval’s strong correlation with Chatbot Arena rankings, making it a reasonable proxy for human judgments (Dubois et al., 2024a). We adopt Ranking Correction, measured by the Spearman rank correlation coefficient, to evaluate the similarity. To generate rankings, we utilize outputs from the open-source models Llama3-70B, Qwen2-72B, Mistral-8$\times$7B, and Qwen1.5-72B as our evaluators. For further implementation details, see Appendix A. We investigate two variants of GED: (w/o ensemble) denoises the preference graphs from different evaluators for the same question, converts each into a ranking, and then ensembles these rankings to produce the final output, while (w/o denoising) directly ensembles the preference graphs to obtain the final ranking without denoising.
Figure 3: Comparison of GED and (w/o ensemble) variants. GED outperforms due to preserving more information by directly ensembling preference graphs, while rank aggregation in the (w/o ensemble) methods leads to performance loss.
@@
The results, presented in Table 2, show that GED outperforms all single-model baselines, highlighting the significant improvement in ranking accuracy achieved by leveraging preference information from multiple evaluators. Moreover, GED surpasses the (w/o ensemble) variant, indicating that generating rankings through graph ensemble first prevents information loss compared to converting individual graphs into rankings. When the ensemble graph is not denoised (w/o denoising), residual noise can adversely affect the final ranking quality. Additionally, our denoising method also enhances results in single-model settings.
Figure 4: Performance comparison of different methods (Random, Longest, ContraSolver, and GED) across multiple benchmarks, including LIMA, Vicuna, Koala, WizardLM and Self-Instruct. The results show GED effectively filters low-quality responses, improving performance and model alignment over baselines.
Table 3: Performance comparison of different methods (Random, Longest, ContraSolver, and GED) on model alignment task across the HH-RLHF benchmark. The results demonstrate the superiority of GED in consistently selecting high-quality responses, leading to improved model performance compared to baseline methods.
Base Model Harmless (base) Helpful (base) Helpful (online) Helpful (rejection) Avg.
Origin 69.67 61.12 65.41 64.06 65.07
Random 69.38 62.87 66.75 65.57 66.14
Longest 69.65 63.54 66.99 66.43 66.65
ContraSolver 69.57 63.61 66.87 66.59 66.66
Llama-2-7B GED 69.71 64.10 67.87 67.01 67.17
Origin 61.59 59.51 65.21 63.17 62.37
Random 59.15 59.61 64.06 62.38 61.30
Longest 61.81 60.53 64.52 63.22 62.52
ContraSolver 61.48 59.85 64.66 63.41 62.85
Mistral-7B GED 61.96 60.71 65.49 63.82 63.50
7 Experiments on Instruct Tuning
Experiment Setup.
In this section, we explore the effects of various data selection methods for model alignment on Llama-2-7B (Touvron et al., 2023) and Mistral-7B (Jiang et al., 2023) through instruct tuning. Specifically, we randomly sampled 5,000 data points from UltraFeedback (Cui et al., 2023) and used Qwen1.5-14B (Yang et al., 2024a) to generate eight responses per data point as instruct data. We then applied four different methods—Random, Longest (Zhao et al., 2024), ContraSolver (using Qwen2-72B as the evaluator) (Zhang et al., 2024b), and our proposed GED, which leverages Llama3-8B, Mistral-7B, and Qwen2-7B as evaluators—to select a subset of these responses for model alignment training. The Origin refers to the performance of the base model without alignment. The models were evaluated on the HH-RLHF (Bai et al., 2022) benchmark, which comprises four subsets: Harmless (base), Helpful (base), Helpful (online), and Helpful (rejection). For evaluation, we employed the same Reward model as in prior work (Song et al., 2024; Yu et al., 2023) to quantify human preference levels. The results are presented in Table 3. For further implementation details, see Appendix A. To ensure a comprehensive assessment, we further evaluated the models—using the Llama-2-7B backbone—on additional benchmarks, including LIMA (Zhou et al., 2023), Vicuna (Chiang et al., 2023), Koala (Vu et al., 2023), WizardLM (Xu et al., 2023), and Self-Instruct (Wang et al., 2022), in accordance with recent studies (Chen et al., 2023b; Zhang et al., 2024a; Hu et al., 2024). The corresponding results are summarized in Figure 4.
Main results.

### v4

@@ 5 Theoretical Analysis
In this section, we provide a theoretical foundation for our method, showing that by modeling preference graphs as random perturbations of a ground truth DAG, PGED can reliably recover the true structure through graph ensemble and denoising with high probability, demonstrating its robustness in handling noisy evaluations. Theoretically, we treat each of our preference graph as a random perturbation of some ground truth DAG $G=(V,A)$. Specifically, we consider a random graph generator $\mathcal{G}(G,\delta_{1},\delta_{2})$ with parameters $\delta_{1},\delta_{2}\in[0,1]$ such that $G_{i}=(V_{i},A_{i})\sim\mathcal{G}(G,\delta_{1},\delta_{2})$ satisfies $V_{i}=V$.
@@
Now, given that $G_{1},\dots,G_{N}\overset{\text{i.i.d.}}{\sim}\mathcal{G}(G,\delta_{1},\delta_{2})$, we will show that to some extent our combination of graph ensemble and graph denoising can indeed provably recover the ground truth DAG $G$. For simplicity, all edges in $G_{1},\dots,G_{N}$ and $G$ are considered equal weighted. Meanwhile, we use $\text{MAS}(\cdot)$ to denote the graph obtained by denoising, which stands for the maximum acyclic subgraph (MAS). Then, we have the following theorem.
Theorem 1
Suppose $G_{1},\dots,G_{N}\overset{\text{i.i.d.}}{\sim}\mathcal{G}(G,\delta_{1},\delta_{2})$ for some ground truth $G=(V,A)$. Let $\widehat{G}$ be the graph ensembled from $G_{1},\dots,G_{N}$ by operations defined in Section 3. Then, as long as $\delta_{1}=0.5-\epsilon$ for some $\epsilon>0$, we have
@@
The full proof is given in Appendix J. From the theorem, we can see that the probability of failure decreases exponentially as the number of samples $N$ increases. Meanwhile, this guarantee only requires $\delta_{1}<0.5$ and does not place restrictions on $\delta_{2}$, which are very mild conditions.
@@
Majority Voting 66.18 29.57 74.77 86.42 11.72 53.73
Weight Majority Voting 66.43 30.12 74.79 86.68 11.84 53.97
PRD 66.53 30.33 74.89 86.75 11.83 54.07
ChatEval 66.51 30.26 74.79 86.85 11.92 54.07
SPA 66.32 30.05 74.82 86.57 11.81 53.91
PGED(w/o denoising) 69.25 30.98 74.29 87.17 12.68 54.87
Multiple evaluator PGED 71.86 32.95 76.57 89.76 13.74 56.98
Table 1: Performance comparison of response selection methods across five benchmarks. PGED consistently outperforms baseline methods, highlighting the effectiveness of graph denoising and multi-evaluator aggregation. Single model denotes directly using the evaluator to answer the question. Single evaluator and Multi evaluator refer to selecting the best response from ten Qwen2-72B-generated candidates using a single or multiple evaluators, respectively.
6 Response Selection for Test-time Scaling
@@
In this section, we evaluate the performance of PGED on five benchmarks: HumanEval (Chen et al. 2021), AlpacaEval (Li et al. 2023b), MATH (Hendrycks et al. 2021), GSM8k (Chen et al. 2021), and GAIA (Mialon et al. 2023). The Qwen2-72B (Yang et al. 2024) model ($\mathcal{M}$) generates ten candidate responses per question, and we assess the effectiveness of different methods in selecting the best response. We evaluate performance using three setups. First, in the single model setting, the baselines include ContraSolver(Zhang, Yin, and Wan 2024), Self-consistency(Wang et al. 2022), and direct evaluation with models (Llama3-8B, Mistral-7B, Qwen2-7B and Qwen2-72B). Additionally, we include a baseline called ListPreference, where instead of pairwise comparisons, all candidate responses are input into Qwen2-72B for selecting the most appropriate response. Then, in the single evaluator setting, individual evaluators (Llama3-8B, Mistral-7B, Qwen2-7B, Qwen2-72B) select the best response from $\mathcal{M}$’s outputs, with and without applying PGED’s graph denoising. We additionally report results for three classical baselines in this setting: Elo (Jiang, Ren, and Lin 2023), Bradley–Terry (BTL) (Bradley and Terry 1952; Lu and Boutilier 2011), and MergeSort (Shyamasundar 1996). Finally, in the multiple evaluators setup, we combine three small evaluators (Llama3-8B, Qwen2-7B, Mistral-7B) to select responses from Qwen2-72B with PGED. We additionally include four aggregation baselines: Majority Voting, Weighted Majority Voting, PRD (Li, Patel, and Du 2023a), ChatEval (Chan et al. 2023a), and SPA (Kadekodi, McTavish, and Ustun 2025). We present the results of PGED and its variant (w/o denoising), which ensembles the preference graphs without the denoising step.
Note that, among the baselines we report: (i) in the single model setting, ContraSolver, ListPreference, and Self-consistency use Qwen2-72B as the base model; (ii) in the single evaluator setting, Elo, Bradley–Terry, and MergeSort use Qwen2-72B as the base model; and (iii) in the multiple evaluators setting, Majority Voting, Weighted Majority Voting, PRD, ChatEval and SPA use Llama3-8B, Mistral-7B, and Qwen2-7B as the base model set. For details on the datasets and baselines, please refer to Appendix A.
Figure 2: Comparison of PGED with GPT-3.5, GPT-4-o-mini, and GPT-4-o on 100 randomly selected tasks. PGED consistently outperforms GPT-3.5 across all tasks and surpasses GPT-4-o-mini on challenging tasks like HumanEval and GSM8k, showcasing the effectiveness of weak evaluator aggregation with graph denoising.
Main results.
Table 1 presents the results of the response selection task across five benchmarks. PGED consistently outperforms all baseline methods, including single model evaluations (single model), direct response selection by individual models (single evaluator) and Multiple evaluator. This demonstrates the strength of aggregating preference evaluators with PGED, particularly when coupled with graph denoising, which enhances response quality by filtering out noise and biases. Furthermore, by combining preference graphs derived from smaller models (Llama3-8B, Mistral-7B, Qwen2-7B), PGED outperforms a much larger evaluator (Qwen2-72B). This underscores the value of ensemble methods in mitigating the limitations of individual evaluators. Then, the denoising process proves to be crucial for improving consistency and overall response quality. The substantial performance gains observed when using PGED with denoising, compared to both the single evaluator setup and the ensemble without denoising, highlight its importance in refining response selection. For Majority Voting, while it improves upon individual evaluators, it still underperforms PGED, highlighting PGED’s ability to capture nuanced evaluation signals and reduce inconsistencies. Additionally, we observed that the ListPreference baseline performed worse than Qwen2-72B as single evaluator, likely due to LLM limitations in handling long-text. Lastly, to further evaluate PGED, we compared its performance with GPT-3.5, GPT-4-o-mini, and GPT-4-o. Due to computational and API cost constraints, we limited the evaluation to 100 data points for each task. As shown in Figure 2, PGED consistently outperformed GPT-3.5 across all tasks and surpassed GPT-4-o-mini on challenging benchmarks like HumanEval and GSM8k. These results highlight the superiority of PGED, particularly in leveraging multi-preference evaluators and graph denoising to outperform individual state-of-the-art models. More discussion on cost is provided in Appendix R.
@@
PGED(w/o ensemble) 62.82 68.44 68.44 69.34 67.34 67.27
PGED(w/o denoising) 64.84 69.23 69.81 75.35 74.37 70.72
Multiple evaluator PGED 66.59 71.14 71.14 77.17 76.46 72.50
7 Model Ranking for Evaluation
Experiment Setup.
[…]
@@ 7 Model Ranking for Evaluation
The results, presented in Table 2, show that PGED outperforms all single-model baselines, highlighting the significant improvement in ranking accuracy achieved by leveraging preference information from multiple evaluators. Moreover, PGED surpasses the (w/o ensemble) variant, indicating that generating rankings through graph ensemble first prevents information loss compared to converting individual graphs into rankings. When the ensemble graph is not denoised (w/o denoising), residual noise can adversely affect the final ranking quality. Additionally, our denoising method also enhances results in single-model settings.
Figure 3: Performance comparison of different methods (Random, Longest, ContraSolver, and PGED) across multiple benchmarks. The results show PGED effectively filters low-quality responses, improving performance and model alignment over baselines.
8 Data Selection for Model Fine-tuning
@@
Base Model Harmless (base) Helpful (base) Helpful (online) Helpful (rej.) Avg.
Origin 69.68 61.65 64.84 63.37 64.89
Random 69.85 63.53 67.00 65.85 66.56
Longest 69.32 64.15 66.56 66.30 66.58
ContraSolver 69.75 63.28 66.71 65.86 66.40
Llama-2-7B
Ours 72.27 65.56 69.44 67.60 68.72
Origin 61.30 59.95 65.21 63.75 62.55
Random 59.85 59.34 63.38 61.87 61.11
Longest 61.66 59.73 64.92 62.92 62.31
ContraSolver 60.87 59.83 64.23 63.26 62.05
Mistral-7B
Ours 64.57 64.48 69.66 67.73 66.61
Table 3: Performance comparison of different methods on model alignment task across the HH-RLHF benchmark.
In this section, we explore the effects of various data selection methods for model alignment on Llama-2-7B (Touvron et al. 2023) and Mistral-7B (Jiang et al. 2023) through instruct tuning. Specifically, we randomly sampled 5,000 data points from UltraFeedback (Cui et al. 2023) and used Qwen1.5-14B (Yang et al. 2024) to generate eight responses per data point as instruct data. We then applied four different methods—Random, Longest (Zhao et al. 2024), ContraSolver (using Qwen2-72B as the evaluator) (Zhang, Yin, and Wan 2024), and our proposed PGED, which leverages Llama3-8B, Mistral-7B, and Qwen2-7B as evaluators—to select a subset of these responses for model alignment training. The Origin refers to the performance of the base model without alignment. The models were evaluated on the HH-RLHF (Bai et al. 2022) benchmark, which comprises four subsets: Harmless (base), Helpful (base), Helpful (online), and Helpful (rejection). For evaluation, we employed the same Reward model as in prior work (Song et al. 2024; Yu et al. 2023) to quantify human preference levels. The results are presented in Table 3. To ensure a comprehensive assessment, we further evaluated the models—using the Llama-2-7B backbone—on additional benchmarks, including AlpacaEval 2.0 (Li et al. 2023b), LIMA (Zhou et al. 2023), Koala (Vu et al. 2023), and Self-Instruct (Wang et al. 2022), in accordance with recent studies (Chen et al. 2023b; Zhang et al. 2024; Hu et al. 2024). The corresponding results are summarized in Figure 3. For details on the datasets and baselines, please refer to Appendix A.